OpenStudy (anonymous):

A surveyor measured a triangular plot of ground and reported the three angles as being 58.2 degrees, 70.0 degrees and 61.1 degrees. How much of an error did the surveyor make? 8.7 3.1 15.3 9.3

OpenStudy (radar):

How many degrees to the angles of a triangle sum up to: Or another way of sayin, How many degrees in a triangle?

OpenStudy (radar):

*to do

OpenStudy (anonymous):

180

OpenStudy (anonymous):

Use the triangle sum theorem. 58.2 + 70 + 61.1 must equal 180 If it doesn't, the amount of error is the difference between the two measures.

OpenStudy (radar):

How many degrees did the surveyors triangle have ?

OpenStudy (anonymous):

9.3 is the difference

OpenStudy (radar):

You will see the answer when you answer that.

OpenStudy (radar):

Yes

OpenStudy (anonymous):

my brain is so fried thank you would you help out with a couple more

OpenStudy (anonymous):

you made that so easy i wish i had an instructor that would make it that easy

OpenStudy (radar):

If I am capable. I will try.

OpenStudy (anonymous):

What is the height of a triangle of area 70 square inches if its base is 20 inches? A. 3.5 inches B. 14 inches C. 7 inches D. 5 inches

OpenStudy (radar):

Do you know the formula for the area of a triangle?

OpenStudy (anonymous):

a2+b2=c2

OpenStudy (anonymous):

my bad its 7

OpenStudy (anonymous):

a=1/2bh

OpenStudy (radar):

If you look it up, it is: A = 1/2 base times the height. That formula is a very important formula for a right triangle. It is saying the hypotenuse (the longest side) squared is equal to the sum of the square of each of the other two sides (a,b) This does not apply here as we don't even know for sure it is a right triangle. The area formula is the one we can use to solve this problem.

OpenStudy (radar):

Hey you are way ahead of me lol. OK lets plug in the values we know: 70 =1/2(20) h The height is the unknown quantity we are attempting to solve.

OpenStudy (radar):

Lets work what we have: 70=10h Can you solve for h?

OpenStudy (radar):

Or we could turn it around and have 10h=70

OpenStudy (radar):

Did you understand where I plugged the base (20) into the equation? The area of 70 ?

OpenStudy (radar):

Ary stuck @BuffaloBill?

OpenStudy (radar):

* are you

OpenStudy (anonymous):

sorry my mouse batteries ran out, yup got it all

OpenStudy (radar):

What did you come up for h the height?

OpenStudy (anonymous):

yes